OpenStudy (anonymous):

A-4 cylinder car typpically gets 21 miles for every gallaon of gas.The odometer reading on the car is 22,050 miles.How many gallons of gas been used since the car was purcesed?Please sho and explain work

OpenStudy (anonymous):

1050 I divided it

OpenStudy (jamesj):

So the car has travelled 22,050 miles. How many miles does it travel with 1 gallon? 21. Hence if it has travelled X miles, it has consumed X/21 gallons. Use that expression.

OpenStudy (anonymous):

i dont no if im rite

OpenStudy (anonymous):

so what is the answer james?

OpenStudy (anonymous):

21 miles / 1 gallon = 22,050 miles / X gallons (it's a ratio, and you need to solve for X) 21x=22050(1) x = 22050/21 x = 1050 gallons of gas

OpenStudy (anonymous):

told yaa